OpenStudy (anonymous):

The male and female guests attending Patricia's party were surveyed to see if they drove to the party or did not drive. The data is displayed in the table below. Male Female Drove to the party 26 34 Did not drive to the party 34 116 If a guest is chosen at random from this group, what are the chances of choosing a guest who is male and did not drive to the party?

OpenStudy (anonymous):

Okay so first we need to find the amount of males from females. and the amount that drove vs. not drive.

OpenStudy (anonymous):

So 26+34 = 60

OpenStudy (anonymous):

60 females 150 males, 60 that drove and 150 that didn't

OpenStudy (anonymous):

26/60 = 13/30 is the percentage of males.

OpenStudy (anonymous):

Hmm. Is your Data correct. Because where do you get 116 people that drove when there's not even that many males and females added together.

OpenStudy (anonymous):

Up there I see 26 males and 34 females which addds up to 60 people.

OpenStudy (anonymous):

there is a total of 116+34 females

OpenStudy (anonymous):

Oh Wow I couln't read the table. Haha. I see it now

OpenStudy (anonymous):

loll oh sorry

OpenStudy (anonymous):

Okay so 60 Males all together and 150 females. Is that right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Alright that gives us 210 total people. Now we need to find the percent that are male so we need to do 60/210 which is 2/7 are male.

OpenStudy (anonymous):

so 28%

OpenStudy (anonymous):

Now we need to find the males that drove and didn't drive. So we do 26+34 = 60

OpenStudy (anonymous):

34/60 which is 17/30

OpenStudy (anonymous):

56%?

OpenStudy (anonymous):

Now we multiply them 2/7 * 17/30

OpenStudy (anonymous):

Let me do it in decimal form so it's easier

OpenStudy (anonymous):

I got about .161

OpenStudy (anonymous):

Which is 16 percent.

OpenStudy (anonymous):

approximately.

OpenStudy (anonymous):

thats awesome!! thanks so much

OpenStudy (anonymous):

it was right!